OpenStudy (anonymous):

Suppose 4x^2+9y^2=36, where x and y are functions of t. If dy/dt=1/3, find dx/dt when x=2 and y=2/3*sqrt(5)

OpenStudy (anonymous):

Diffferentiate implitly, you get: 8x dx/dt + 18y dy/dt = 0 Now, substitute, you get 8(2) dx/dt + 18(1/3)(2/3) = 0 16 dx/dt + 4 = 0 16 dx/dt = -4 dx/dt = -4/16 or -1/4

OpenStudy (anonymous):

implicitly*

OpenStudy (anonymous):

The 2/3 actually \[2/3 \sqrt{5}\] so it would be \[-4\sqrt{5}/16\] right?

OpenStudy (anonymous):

Yes..I didnt see that 3sqrt(5)...you're absolutely correct.

OpenStudy (anonymous):

okay thank you!!